In a survey, three out of four students said that courts show too much concern for criminals. Find the probability that of seventy randomly selected students


Exactly sixty-eight feel that courts show too much concern.
Answer #1

Let A be the event that a student thinks that courts show too much concern for criminals .

We are given with p = P[A] = 3/4

Note that A can be thought of as a "success" in Binomial model .

So if X be the number of students feeling that courts show too much concern then

X~ Bin( n , p=3/4) , x = 0,1,2,....,n

Here n = 70

To find P[X=68]